OpenStudy (anonymous):

5. Given that the variable t represents the temperature at 3 p.m., which equation would best model the word problem below? The temperature at the summit of Mt Mitchell dropped 23°F between 3 p.m. and 9 p.m. If the temperature at 9 p.m. was –10°F, what was the temperature at 3 p.m.? (Points : 1) t – 23 = –10 10 – 23 = t 10 + 23 = t 23 – t = –10

OpenStudy (anonymous):

help? @greenlegodude57

OpenStudy (anonymous):

Hold on I'm thinking . . .

OpenStudy (anonymous):

I think the answer is A.

OpenStudy (anonymous):

Because it was -10F at 9:00 so -10 + 23 = 13. And 13 + 23 = -10.

OpenStudy (anonymous):

But I might be wrong let me double check.

OpenStudy (anonymous):

My bad I think it's D. Before when I thought it was A. I added but It said it dropped so I had to subtract. So the answer is D.

OpenStudy (anonymous):

Plug in "t" on the equation for D. 23 - 33 = -10.

OpenStudy (anonymous):

It's correct.

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

You got it right?

OpenStudy (anonymous):

yes